OpenStudy (anonymous):

(x^2 + 4^x – 3)(2x^2 + x + 6) This Equation is Multiplying a monomial by a polynomial.

OpenStudy (anonymous):

Help Plx!

OpenStudy (anonymous):

I give medals ! (:

OpenStudy (anonymous):

What do we have to do?

OpenStudy (anonymous):

Distribute the #1 equation into #2

OpenStudy (amistre64):

just multiply .... its the same thing you learned back in like the 2nd grade

OpenStudy (amistre64):

x^2 + 4^x – 3 2x^2 + x + 6 --------------- 6x^2 +24^x - 18 then move to the next row

OpenStudy (amistre64):

6*4^x might be more better tho :)

OpenStudy (anonymous):

I think you have the wrong idea .... I need to multiply x^2 times 2x^2 ....

OpenStudy (amistre64):

x^2 + 4^x – 3 2x^2 + x + 6 ------------------------ 6x^2 +6.4^x - 18 x^3 +x.4^x -3x helps to line up corresponding parts tho

OpenStudy (amistre64):

i got the right idea .....

OpenStudy (amistre64):

its just not the "method" that you are proscribing even tho it is exactly the same ....

OpenStudy (anonymous):

Maybe this can help ^

OpenStudy (amistre64):

x^2 + 4^x – 3 2x^2 + x + 6 ------------------------ 6x^2 +6.4^x - 18 x^3 +x.4^x -3x 2x^4 -6x^2+2x^2.4^x ------------------------------- 2x^4+x^3+0 +(3x+6)4^x -3x - 18

OpenStudy (amistre64):

that does help :) you posted 4^x by mistake

OpenStudy (anonymous):

OOOOO !! Sorry

OpenStudy (amistre64):

but the same thing applies; multiply it out ... with polys I like to start on the left tho x^2 + 4x – 3 2x^2 + x + 6 ------------- 2x^4 +8x^3 -6x^2 +x^3 +4x^2 -3x +6x^2 +24x-18 -------------------------- 2x^4 +9x^3 +4x^2 +21x -18

OpenStudy (anonymous):

I need something like this .....

OpenStudy (amistre64):

thats teaching a proscribed method and not really asking for results. but it looks to mimic my first row

OpenStudy (anonymous):

But his equation is another example. but i need this equation like it.... But i cannt get x^2 * 2x^2

OpenStudy (amistre64):

x^2 = x.x 2x^2 = 2.x.x 2x^2 * x^2 = 2.x.x.x.x = 2x^4

OpenStudy (amistre64):

im not sure what the method is; it seems to be material specific tho

OpenStudy (anonymous):

ok thanks this is what I am looking for..... :)